OpenStudy (jskhupmang):

The radius r of a circle is increasing at a rate of 6 centimeters per minute. a. Find the rate of change of the area when r = 10 centimeters. b. Find the rate of change of the area when r = 28 centimeters.

OpenStudy (jskhupmang):

The answers should be in cm^2/min.

OpenStudy (mrnood):

OK - you want to use calculus?

OpenStudy (jskhupmang):

yes

OpenStudy (mrnood):

so write the equation for area of a circle

OpenStudy (batman*):

this is calculus using derivatives right?

OpenStudy (jskhupmang):

yes

OpenStudy (mrnood):

so write the equation for area of a circle

OpenStudy (jskhupmang):

I think its area=r^2Pi

OpenStudy (mrnood):

correct so what is dA/dr ( the derivative)

OpenStudy (jskhupmang):

\[\frac{ dA }{ dt }=\frac{ d }{ dt }\left[ r^2 \pi \right]\]

OpenStudy (jskhupmang):

right?

OpenStudy (jskhupmang):

but im not sure how to do Product rule for r^2 pi

OpenStudy (batman*):

You will actually get da/dt = pi(2r * dr/dt). This is using both the chain rule and power rule

OpenStudy (xapproachesinfinity):

What Batman is saying is that r is a function of t as well so we require to do chain rule

OpenStudy (batman*):

yes. very much so.

OpenStudy (xapproachesinfinity):

Any further question?

OpenStudy (batman*):

A good way too look at this and how to solve it is this video: https://www.youtube.com/watch?v=KFMHUuCnOsU&list=PLsJ7E_UWMp2wTSq8C0wpJVuwQWbSaqHxZ

OpenStudy (jskhupmang):

so do I just plug in the radius?

OpenStudy (xapproachesinfinity):

Batman pretty much did all, you just need to plug in the values

OpenStudy (batman*):

you have to insert two values. you need to plug in dr/dt and the r. We know what both values are.

OpenStudy (xapproachesinfinity):

notice dr/dt =6

OpenStudy (jskhupmang):

da/dt=pi(2(10cm)*(dr/dt) so what is dr/dt ?

OpenStudy (jskhupmang):

ok now it makes sense :)

OpenStudy (xapproachesinfinity):

Ok good:)

OpenStudy (batman*):

Remember what dr/dt actually is. its a slope, or a rate of change. :)

OpenStudy (batman*):

I'm actually working on a similiar problem right now. haha

OpenStudy (jskhupmang):

THanks i was thinking about what would be dr/dt. Thank you all!

OpenStudy (jskhupmang):

thanks @BATMAN*